OpenStudy (anonymous):

the center of a circle represented by the equation (x-2)^2+(y+3)^2 =100 is located in Quadrant?

OpenStudy (jdoe0001):

http://www.mathwarehouse.com/geometry/circle/images/equation-of-circle/general-formula-equation-of-circle.png \(\begin{array}{llll} (x-2)^2+(y+3)^2 =100\implies (x-&2)^2+(y-(&-3))^2 =&100\\ &\uparrow&\uparrow\\\ &h&k&r \end{array}\\ \quad \\\bf center\implies (h,k)\)

OpenStudy (jdoe0001):

find that point, see on what Quadrant it's at